What is the center radius form of the equation of the circle

1
Expert's answer
2022-01-24T18:48:15-0500

"x^2+y^2-2x+4y-1=0\\
x^2+y^2-2x+4y=1"

Add the square of the half of the coefficient of "x" and "y" to both sides.

"x^2+y^2-2x+(-1)^2+4y+(2)^2=1+(-1)^2+(2)^2\\
x^2-2x+(-1)^2+y^2+4y+(2)^2=1+1+4\\
(x-1)^2+(y+2)^2=6"

Compare this to the standard equation of a circle which is:


"(x-a)^2+(y-b)^2=r^2"

This implies that,


"a=1, b=-2, r=\sqrt6"